Suppose we want to form three-digit numbers using the set of digits {0,1,2,3,4,5}. How many of these numbers are multiples of 5?
Answer by edjones(8007)

nPr=Permutation of n things taken r at a time.=n!/(n-r)!
6!/3! = 6*5*4 = 120 arrangements
2/6 = 1/3 of these end in 5 or 0.
120 * 1/3 = 40 of these numbers are multiples of 5.
.
Ed